# Equitable Allocation for Mixtures of Goods and Chores

### v2

@@ Appendix B Omitted Details of Section 5
We will show that the allocation $A$ as returned by the Algorithm 3 is EQ1$+$PO. Notice that $A$ allocates all items from $O^{\pm}\cup O^{+}$ non-wastefully. Also, $o\in O^{-}$ is allocated to an agent who values it at 00, otherwise, if both the agents value it at $-1$, then PO is satisfied irrespective of which agent ends up receiving that item. Therefore, $A$ satisfies PO. We now argue that it also satisfies EQ1. Suppose $i$ is a poor agent by the end of Step $6$. Say, the utility of $i$ at this point is $k_{1}\cdot c$ and that of agent $j$ is $k_{2}\cdot c$. Since $i$ is the poor agent, therefore, $k_{1}<k_{2}$. Then, there must be at least $k_{2}-k_{1}$ items in $O^{+}$ valued at $\{1,0\}$ by $\{i,j\}$ (by type-normalization). By construction, all these $k_{2}-k_{1}$ items are allocated to $i$ under $A$. This compensates for the inequity experienced by $i$ so far. Now the allocation is extended by iteratively allocating a poor agent its most valuable item, which maintains EQ1 till $O^{+}$ is exhausted. For the items in $O^{-}$ which are valued at $-1$ by both agents, the rich agent gets that item, which thereby maintains EQ1. For the remaining chores in $O^{-}$, at least one of the agents values them at 00 and ends up receiving the same. This does not violate EQ1 as it does not change the utility of the agents. This settles the claim. ∎
Appendix C Omitted Details of Section 6
@@
(continued) We argue here that every entry in the DP table is indeed computed correctly. To that end, we do an induction on the number of items allocated. For the base case, when one item is allocated to say, agent $i$, then only $i$ derives the value of $v_{i}(o_{1})$, and the rest of the agents get a value 00. Depending on whether $o_{1}$ is a good or a chore for agent $i$, either $g_{i}=o_{1}$ or $c_{i}=o_{1}$, and everything else is $\emptyset$. This is correctly captured in the base case. By induction hypothesis, suppose all the table entries that allocate the first $k-1$ items are computed correctly. Consider the allocation of $k^{th}$ item as captured by the table entry $(k,\mathbf{v},\mathbf{g},\mathbf{c})$.
First consider the case when $g_{i}\neq o_{k}$ and $c_{i}\neq o_{k}\leavevmode\nobreak\ \forall i\in[n]$ in the entry $(k,\mathbf{v},\mathbf{g},\mathbf{c})$. Suppose RHS of Section 6 is True. Suppose $o_{k}$ is a good for some agent $i$ such that $v_{i}(g_{i})\leq v_{i}(o_{k})$ (the case of chores can be argued similarly) and RHS of Section 6 is True for the index $i$. That is, $(k-1,v_{1},\ldots,v_{i}-v_{i}(o_{k}),\ldots v_{n};g_{1},\ldots g_{i},\ldots g_{n};c_{1},\ldots c_{i},\ldots c_{n})$ is True. It means that there is allocation of $k-1$ items such that everyone gets a utility of $(v_{1}\ldots v_{i}-v_{i}{(o_{k})},\ldots v_{n})$ and $(g_{1}\ldots g_{n})$ and $(c_{1},\ldots c_{n})$ are the least valued goods and the least disliked chores in the respective bundles. Then allocating $o_{k}$ to $i^{th}$ agent gives an allocation with utilities $(v_{1},\ldots v_{i},\ldots v_{n})$ such that the set of the least valued goods and least disliked chores remain the same for all the agents (because $v_{i}(g_{i})\leq v_{i}(o_{k})$). Therefore, LHS of Section 6 is True.
On the other hand, suppose the LHS of Section 6 is True. Suppose $o_{k}$ is a good and belongs to agent $i^{\prime}s$ bundle. Then, since $g_{i}\neq o_{k}$, it means that $g_{i}$ is the least valued item in agent $i^{\prime}s$ bundle. Consider the allocation after removing $o_{k}$ from $i^{\prime}s$ bundle. Then, it corresponds to an allocation of $k-1$ items such that each agent gets a utility of $(v_{1},\ldots v_{i}-v_{i}{(o_{k})},\ldots v_{n})$ such that $(g_{1},\ldots g_{n})$ and $(c_{1},\ldots c_{n})$ are the least valued goods and the least valued chores in the respective bundles. Therefore, $(k,v_{1},\ldots v_{i}-v_{i}{(o_{k})},\ldots v_{n},g_{1},\ldots g_{n},c_{1},\ldots c_{n})$ is True and hence RHS of Section 6 is True.
@@ Appendix D EF+EQ+PO Allocations
Suppose $A$ is an EQ+PO allocation for the given instance. EQ implies that we have $v_{i}(A_{i})=v_{i}(A_{j})=k\cdot c$ and PO ensures that if an agent receives an item $o$ with value $-1$, then everyone else values $o$ at $-1$. (Else, if there is an agent $i$ such that $v_{i}(o)=0\leavevmode\nobreak\ or\leavevmode\nobreak\ c$, then allocating $o$ to $i$ is a Pareto improvement). Likewise, if an agent receives an item that it values at 00, then everyone else values that item at either 00 or $-1$, again for the same reason. Now suppose $A$ is not EF. Then, there is a pair of agents $i$ and $j$ such that $v_{i}(A_{i})=k\cdot c$ but $v_{i}(A_{j})>k\cdot c=v_{j}(A_{j})$. This implies that there is an item $o$ in $j^{\prime}s$ bundle that is valued at 00 (or $-1$) by $j$ but valued at $1$ (or 00) by $i$. Allocating $o$ to $i$ is a Pareto improvement, which is a contradiction. Therefore, $A$ is EF. ∎
@@
Corollary D.3.
Deciding whether a type-normalized instance admits an allocation that is simultaneously EF+EQ1$+$PO or EF1+EQ1$+$PO is (strongly) NP-hard.
[…]
Corollary D.4.
For $\{1,0,-1\}$ valuations, an EF1+EQ1$+$PO allocation can be computed in polynomial time, whenever such an allocation exists.

### v3

@@ Appendix C Omitted Details of Section 5
We will show that the allocation $A$ as returned by the Algorithm 3 is EQ1++PO. Notice that $A$ allocates all items from $O^{\pm}\cup O^{+}$ non-wastefully. Also, $o\in O^{-}$ is allocated to an agent who values it at 0, otherwise, if both the agents value it at 1-1, then PO is satisfied irrespective of which agent ends up receiving that item. Therefore, $A$ satisfies PO. We now argue that it also satisfies EQ1. Suppose $i$ is a poor agent by the end of Step $6$. Say, the utility of $i$ at this point is $k_{1}\cdot c$ and that of agent $j$ is $k_{2}\cdot c$. Since $i$ is the poor agent, therefore, $k_{1}<k_{2}$. Then, there must be at least $k_{2}-k_{1}$ items in $O^{+}$ valued at $\{1,0\}$ by $\{i,j\}$ (by type-normalization). By construction, all these $k_{2}-k_{1}$ items are allocated to $i$ under $A$. This compensates for the inequity experienced by $i$ so far. Now the allocation is extended by iteratively allocating a poor agent its most valuable item, which maintains EQ1 till $O^{+}$ is exhausted. For the items in $O^{-}$ which are valued at 1-1 by both agents, the rich agent gets that item, which thereby maintains EQ1. For the remaining chores in $O^{-}$, at least one of the agents values them at 0 and ends up receiving the same. This does not violate EQ1 as it does not change the utility of the agents. This settles the claim. ∎
Appendix D Omitted Details of Section 6
@@
(continued) We argue here that every entry in the DP table is indeed computed correctly. To that end, we do an induction on the number of items allocated. For the base case, when one item is allocated to say, agent $i$, then only $i$ derives the value of $v_{i}(o_{1})$, and the rest of the agents get a value 0. Depending on whether $o_{1}$ is a good or a chore for agent $i$, either $g_{i}=o_{1}$ or $c_{i}=o_{1}$, and everything else is $\emptyset$. This is correctly captured in the base case. By induction hypothesis, suppose all the table entries that allocate the first k1k-1 items are computed correctly. Consider the allocation of $k^{th}$ item as captured by the table entry $(k,\mathbf{v},\mathbf{g},\mathbf{c})$.
First consider the case when $g_{i}\neq o_{k}$ and $c_{i}\neq o_{k}\leavevmode\nobreak\ \forall i\in[n]$ in the entry $(k,\mathbf{v},\mathbf{g},\mathbf{c})$. Suppose RHS of Section 6 is True. Suppose $o_{k}$ is a good for some agent $i$ such that $v_{i}(g_{i})\leq v_{i}(o_{k})$ (the case of chores can be argued similarly) and RHS of Section 6 is True for the index $i$. That is, $(k-1,v_{1},\ldots,v_{i}-v_{i}(o_{k}),\ldots v_{n};g_{1},\ldots g_{i},\ldots g_{n};c_{1},\ldots c_{i},\ldots c_{n})$ is True. It means that there is allocation of k1k-1 items such that everyone gets a utility of $(v_{1}\ldots v_{i}-v_{i}{(o_{k})},\ldots v_{n})$ and $(g_{1}\ldots g_{n})$ and $(c_{1},\ldots c_{n})$ are the least valued goods and the least disliked chores in the respective bundles. Then allocating $o_{k}$ to $i^{th}$ agent gives an allocation with utilities $(v_{1},\ldots v_{i},\ldots v_{n})$ such that the set of the least valued goods and least disliked chores remain the same for all the agents (because $v_{i}(g_{i})\leq v_{i}(o_{k})$). Therefore, LHS of Section 6 is True.
On the other hand, suppose the LHS of Section 6 is True. Suppose $o_{k}$ is a good and belongs to agent $i^{\prime}s$ bundle. Then, since $g_{i}\neq o_{k}$, it means that $g_{i}$ is the least valued item in agent $i^{\prime}s$ bundle. Consider the allocation after removing $o_{k}$ from $i^{\prime}s$ bundle. Then, it corresponds to an allocation of k1k-1 items such that each agent gets a utility of $(v_{1},\ldots v_{i}-v_{i}{(o_{k})},\ldots v_{n})$ such that $(g_{1},\ldots g_{n})$ and $(c_{1},\ldots c_{n})$ are the least valued goods and the least valued chores in the respective bundles. Therefore, $(k,v_{1},\ldots v_{i}-v_{i}{(o_{k})},\ldots v_{n},g_{1},\ldots g_{n},c_{1},\ldots c_{n})$ is True and hence RHS of Section 6 is True.
@@ Appendix E EF+EQ+PO Allocations
Suppose $A$ is an EQ+PO allocation for the given instance. EQ implies that we have $v_{i}(A_{i})=v_{i}(A_{j})=k\cdot c$ and PO ensures that if an agent receives an item $o$ with value 1-1, then everyone else values $o$ at 1-1. (Else, if there is an agent $i$ such that $v_{i}(o)=0\leavevmode\nobreak\ or\leavevmode\nobreak\ c$, then allocating $o$ to $i$ is a Pareto improvement). Likewise, if an agent receives an item that it values at 0, then everyone else values that item at either 0 or 1-1, again for the same reason. Now suppose $A$ is not EF. Then, there is a pair of agents $i$ and $j$ such that $v_{i}(A_{i})=k\cdot c$ but $v_{i}(A_{j})>k\cdot c=v_{j}(A_{j})$. This implies that there is an item $o$ in $j^{\prime}s$ bundle that is valued at 0 (or 1-1) by $j$ but valued at $1$ (or 0) by $i$. Allocating $o$ to $i$ is a Pareto improvement, which is a contradiction. Therefore, $A$ is EF. ∎
@@
Corollary E.3.
Deciding whether a type-normalized instance admits an allocation that is simultaneously EF+EQ1++PO or EF1+EQ1++PO is (strongly) NP-hard.
Notice that for $\{1,0,-1\}$ valuations, a Pareto optimal allocation is EF1 if and only if it is EFX, and is EQ1 if and only if it is EQX. Therefore, the above two results hold for all combinations of X + Y + PO, where X $\in$ {EFX, EF1} and Y $\in$ {EQX, EQ1}. The allocation constructed in Theorem 5.3 can be easily verified to be EF1, thereby confirming the following result.
Corollary E.4.
For $\{1,0,-1\}$ valuations, an EF1+EQ1++PO allocation can be computed in polynomial time, whenever such an allocation exists.